# SpecFuse: Ensembling Large Language Models via Next-Segment Prediction

### v1

@@ 3.2 Main Results
Llama-3-8B 25.1622 9.7688 23.3102 3.5669 -2.9837 69.9865 9.0850
Glm-4-9B 25.8456 10.2618 23.8950 3.4774 -2.9608 70.5125 8.7993
Qwen2-7B 26.6179 10.8107$\ast$ 24.4886 3.8603 -2.9380 71.4384 8.1443
Gemma-2-9B 25.3130 10.0080 23.5932 4.1933 -2.9282$\ast$ 71.5234 8.6027
Mistral-7B 27.7450$\ast$ 10.7520 25.5678$\ast$ 4.8154$\ast$ -2.9368 71.8773$\ast$ 7.5093$\ast$
@@
Llama-3-70B 26.7744 10.8736 24.5639 4.0981 -2.8376 70.9799 5.2153
Qwen2-72B 27.2580 11.2312 25.1139 4.2896 -2.7601 71.7302 4.1950
Mixtral-8x7B 29.0371 12.2504 26.7546 4.0820 -2.8131 72.1949 3.9461
Ensemble Base LLMs
GF (Qwen2) 23.0829 8.9201 21.2768 3.1881 -2.9513 69.7043 9.6143
GF (Gemma-2) 21.8077 7.6626 20.0847 3.0041 -3.0178 68.1968 9.7543
GF (Mistral) 24.9248 9.5800 22.9664 3.9242 -2.9263 70.3801 8.2710
MBR 27.1221 10.4025 25.3322 4.5642 -2.8912 71.6312 7.5003
PairRank 28.2055 10.8611 25.9361 4.9900 -2.8637 72.0871 6.7073
@@
SpecFuse (w/o ET) 30.8566 14.0015 28.5648 5.5113 -2.8801 72.8901 3.8227
Table 1: Performance on the English Open-Domain Instruction-Response benchmark, with the best result for each metric highlighted in bold and an $\ast$ indicating the highest result among base LLMs. The upward arrow indicates that a higher value for the metric is better, while the downward arrow indicates that a lower value is better. All ensemble methods in the table integrate all base LLMs, with GF (Qwen2) using the outputs of the other base LLMs as context to generate the fused result through Qwen2-7B.
@@
Gemma-2-9B 29.1486 7.6523 18.3456 3.3647 -4.2845$\ast$ 68.7312 8.5471
Qwen2-7B 29.9296 8.0901 20.0345 3.6181 -4.3271 69.9989 6.5122
Mistral-7B 30.9890$\ast$ 8.6498 20.6568$\ast$ 4.4205 -4.4800 70.1000 6.6230
Glm-4-9B 30.8761 8.7092$\ast$ 20.4193 4.4674$\ast$ -4.3018 70.2452$\ast$ 5.2068$\ast$
@@
Llama-3-70B 27.7816 7.0486 20.2227 4.1399 -4.5517 68.5211 7.4415
Qwen2-72B 31.4356 8.9688 22.4781 4.8838 -4.3368 70.6480 3.4485
Ensemble Base LLMs
GF (Qwen2) 28.6936 7.8675 18.9339 3.3169 -4.4105 69.8126 8.3508
GF (Mistral) 30.2933 8.1220 20.3324 3.8817 -4.5356 70.0437 7.1736
GF (Glm-4) 30.2643 8.6996 20.5103 4.2722 -4.3316 70.2348 5.5864
MBR 30.9335 8.7132 20.6322 4.3149 -4.3060 70.2266 4.9921
@@
Open-Domain Instruction-Response Tasks. We evaluate the performance of our ensemble framework in responding to user queries on both English and Chinese benchmarks and compare it with single LLMs and other ensemble methods. The test results on the English benchmark are shown in Table 1. The experimental results demonstrate that by integrating base LLMs, SpecFuse surpasses all base LLMs and previous ensemble methods across all metrics, with an average improvement of more than $3$ points in the Rouge (n) scores, while also achieving the highest overall ranking in the GPT4-Rank metric, which the responses of all models in the table using GPT-4222https://openai.com/index/gpt-4/. Compared to large models with over $70$B parameters, our method is competitive across most metrics. This suggests that in open-domain scenarios with uneven instruction difficulty, SpecFuse provides more stable output by integrating the advantage of multiple base LLMs, achieving response quality comparable to larger models while significantly lowering deployment costs and complexity. Previous generation fusion methods, such as GF (Glm-4), exhibit poor integration performance without additional training, and in some metrics, their performance is even worse than that of the individual models. Furthermore, as shown in Table 2, the results on the Chinese benchmark are similar to those on the English benchmark, with SpecFuse outperforming previous ensemble methods and base LLMs across all metrics, demonstrating that its effectiveness is not constrained by language and highlighting its strong generalization ability.
@@
Commonsense, Reasoning, and Instruction-Following Tasks. To investigate the performance of SpecFuse on commonsense, reasoning, and instruction-following tasks, we select three base LLMs with different task specializations and conduct experiments on four benchmarks. As shown in Table 3, SpecFuse (All) outperforms previous ensemble methods across all four benchmarks and is not constrained by the task format. Additionally, by integrating different combinations of base LLMs with SpecFuse, we observe the following: (1) When the performance gap between base LLMs is not very large, SpecFuse delivers the most significant overall improvement, with SpecFuse (Qwen2+GLM4) achieving gains of $+2.5$ on MMLU and $+2.4$ on ARC-C by leveraging the respective strengths of Qwen2-7B and GLM-4-9B. (2) When the performance gap is large, integrating three LLMs yields more stable results compared to two. For instance, in the IFEVAL benchmark, Qwen2-7B lags $20$ points behind Gemma-2-9B on average, and after integration, the performance improves by $15$ points compared to Qwen2-7B but decreases by $5$ points compared to Gemma-2-9B. Adding GLM-4-9B to the ensemble brings the performance becomes roughly the same as Gemma-2-9B, illustrating that relying on a single strong model in an ensemble system is insufficient. Frequent updates or the addition of new models are necessary, as the rapid evolution of large models quickly makes previous SOTA LLMs obsolete, showcasing the plug-and-play advantage of our framework.
4 Analysis
In this section, we first conduct ablation studies to analyze the significance of the model exit mechanism in our approach, followed by an analysis of the maximum candidate segment length, the number of base LLMs, and latency. Additionally, the case study is described in Appendix C.
Ablation On Model Exit Mechanism. We conduct an ablation study of the Model Exit (ET) mechanism on the Instruction-Response English benchmark test set, using five base LLMs. As shown in Table 4, SpecFuse ($\tau=1$) results in the fewest model invocations but suffers from significant performance loss. This occurs because, as iterations increase, the cumulative model scores grow larger, causing the softmax function to produce a sharper distribution. Consequently, models with slightly lower scores are prematurely eliminated from the response, which negatively impacts the overall performance of the ensemble. SpecFuse ($\tau=\sqrt{T}$) sets the temperature coefficient to $\sqrt{T}$, which makes the softmax function overly smooth. As a result, it takes many iterations to accumulate substantial score differences between models before the lower-scoring models exit the current response, leading to delayed exits and an excessive number of model invocations overall. The ET mechanism, with dynamic temperature scaling, adjusts the temperature coefficient based on the distribution of the best candidate from previous rounds, ensuring timely model exits. As shown in Figure 2, the proportion of best candidate generations for each LLM changes only slightly with or without ET, indicating that the mechanism primarily eliminates LLMs with low selection probability, minimizing its impact on overall performance while reducing the number of base LLMs invocations.
Model Bert-S BLEU RougeL AMIR
SpecFuse (w/o ET) 72.8901 5.5113 28.5648 5.0000
SpecFuse ($\tau=1$) 71.7362 4.2908 25.1196 2.0268
SpecFuse ($\tau=\sqrt{T}$) 72.8105 5.3863 28.2746 4.1996
SpecFuse 72.8354 5.2799 28.3507 2.4168
Table 4: Ablation study of the Model Exit mechanism. SpecFuse ($\tau=1$) indicates that the Model Exit Mechanism is used but without dynamic temperature scaling, with the temperature fixed at 1, and ($\tau=\sqrt{T}$) indicates the temperature is fixed at $\sqrt{T}$, where T refers to the current iteration of SpecFuse. AMIR refers to the average number of models invoked per iteration.
Analysis of the Maximum Length of Candidate Segments. To explore the impact of different maximum generation lengths of candidate segments on the performance of the SpecFuse framework, we conducted tests on the English open-domain IR development set. As shown in Figure 3, the BertScore initially rises with increasing maximum length, reaching its highest point at a length of $10$, after which it begins to decline. When the candidate segment length is too short, it contains insufficient information, affecting the verification component’s judgment and making it difficult for models to effectively inspire one another. On the other hand, if the length is too long, the reduced frequency of cross-model interaction leads to less effective knowledge fusion, ultimately diminishing the quality of the final output. Additionally, we observe that as the candidate segment length increases, the first-token latency also increases, but since the total output length does not change significantly, the overall number of system iterations decreases, resulting in more tokens generated per second on average. Based on these observations, we conclude that setting the maximum candidate segment length to $10$ during model inference provides the best generation quality, while starting with a shorter length allows users to receive quicker feedback.
Figure 3: The variation trends of SpecFuse’s BertScore, first-token latency, and tokens generated per second as the maximum generation length of each candidate segment changes.
Analysis of the Number of Base LLMs. We test the variation in SpecFuse’s performance as the number of base LLMs increases on the test set of the Open-Domain Instruction-Response English benchmark. As shown in Figure 5, the performance of SpecFuse consistently improves as the number of integrated base LLMs increases. When stronger models are introduced, the improvements are substantial, while adding weaker models results in moderate enhancements. This reveals that even weaker models contribute to overall system performance by integrating their strengths, highlighting the advantage of our framework in seamlessly incorporating new LLMs without the need for any training or adaptation.
@@ 4 Analysis
Figure 4: The variation in SpecFuse’s RougeL score as the number of base LLMs increases.
Model FTL/s PTL/s
Llama-3-8B 0.3491 0.0231
Glm-4-9B 0.5500 0.0362
Qwen2-7B 0.3551 0.0258
Gemma-2-9B 0.3872 0.0497
Mistral-7B 0.3720 0.0278
GF (Qwen2) 5.3347 0.0801
GF (Gemma-2) 4.0994 0.0924
GF (Mistral) 5.3890 0.0824
MBR 7.1975 0.0720
SpecFuse 1.2883 0.0765
Figure 5: Results of the inference latency comparison experiment. FTL refers to first-token latency, and PTL refers to the per token latency. The generated length is fixed at $100$ tokens.
Analysis of Latency. Since the time users wait for the system to generate the first token (FTL), as well as the average time per token (PTL), significantly affects the user experience, we conduct experiments on the Open-Domain IR English benchmark to compare the first-token latency and the average latency per generated token with other methods. We select $200$ queries from the development set and instruct the models to generate responses of $100$ tokens each. The maximum length of the candidate segments is set to $10$. We can see from Table 5 that in previous ensemble methods, the first-token latency remains around $5$ seconds for a response limited to $100$ tokens, and this time doubles as the response length increases to $200$ tokens, significantly reducing the user experience due to the prolonged waiting time. Compared to them, SpecFuse reduces the first-token latency by $3$ to $6$ times and can further decrease it by adjusting the maximum length of the candidate segments generated in the first round. For PTL, SpecFuse also consumes less time compared to generation fusion methods, generating more tokens per second. Additionally, since it takes approximately $0.2$ seconds for a human to read each token, these models can meet the reading requirements of users. However, the acceptable FTL for users is typically $1$-$2$ seconds, and previous methods frequently encounter the issue of excessively high FTL during online user interactions. In contrast, our model meets this requirement while delivering better ensemble performance.
@@ 5 Related Work
Post-hoc ensemble. The basic idea of the post-hoc ensemble (Jiang et al., 2023b; Lv et al., 2024a) method is that all base large models first generate complete responses independently for a given question, and then these responses are fused to produce the final ensemble response. Jiang et al. (2023b) trained a reward model, PairRank, to compare pairs of candidate results generated by multiple LLMs, selecting the highest-quality candidate as the ensemble output. As these selection-based methods may restrict the ability to fully utilize each candidate’s strengths, Jiang et al. (2023b) proposed LLM-Blender, which first ranks the candidates using PairRank and then uses a trained fusion model to take the top few candidates as context to generate a fused output. Subsequently, Lv et al. (2024b) proposed the URG ensemble method, an end-to-end framework that first ranks responses, selects the top few candidates as context, and then generates a new response.
Pre-selection ensemble. The pre-selection ensemble methods (Lu et al., 2023; Wang et al., 2024; Shnitzer et al., 2023) test base models on a dataset, maps each question to its best-performing model, and then train a routing model using this mapping data to classify future questions to the most proficient model. Lu et al. (2023) trained a reward model that scores the given query and routes it to the highest-scoring model, inferring from that model alone. Similarly, Wang et al. (2024) trained a model called Frugal using expert LLM outputs on training data, and during inference, it obtains encoded vectors from all LLMs for a given query, which are then fed into Frugal to select one expert LLM that produces the final prediction.
However, these two methods focus on either selecting a single LLM’s response or merging the complete responses from LLMs, resulting in overlooking the potential for mutual inspiration among LLMs to collaboratively generate higher-quality responses during inference. Additionally, due to the need for training an additional fusion or routing model on specific datasets, these methods tend to struggle with poor generalization when handling open-domain queries from users. In contrast, our SpecFuse framework generates fused results by iteratively producing the next segment through collaboration among LLMs, which supports easy plug-and-play integration of LLMs without requiring any training, thereby avoiding generalization issues.

### v2

@@ 3.2 Main Results
Llama-3-8B AI@Meta (2024) 25.1622 9.7688 23.3102 3.5669 -2.9837 69.9865 9.6857
Glm-4-9B GLM et al. (2024) 25.8456 10.2618 23.8950 3.4774 -2.9608 70.5125 9.3630
Qwen2-7B Yang et al. (2024) 26.6179 10.8107$\ast$ 24.4886 3.8603 -2.9380 71.4384 8.5854
Gemma-2-9B Gemma et al. (2024) 25.3130 10.0080 23.5932 4.1933 -2.9282$\ast$ 71.5234 9.1670
Mistral-7B-v0.3 Jiang et al. (2023a) 27.7450$\ast$ 10.7520 25.5678$\ast$ 4.8154$\ast$ -2.9368 71.8773$\ast$ 7.9417$\ast$
@@
Llama-3-70B AI@Meta (2024) 26.7744 10.8736 24.5639 4.0981 -2.8376 70.9799 5.4910
Qwen2-72B Yang et al. (2024) 27.2580 11.2312 25.1139 4.2896 -2.7601 71.7302 4.3990
Mixtral-8x7B Jiang et al. (2023a) 29.0371 12.2504 26.7546 4.0820 -2.8131 72.1949 4.2977
Baselines of Ensemble Method
GF (Qwen2) Jiang et al. (2023b) 23.0829 8.9201 21.2768 3.1881 -2.9513 69.7043 10.1847
GF (Gemma-2) Jiang et al. (2023b) 21.8077 7.6626 20.0847 3.0041 -3.0178 68.1968 10.3030
GF (Mistral) Jiang et al. (2023b) 24.9248 9.5800 22.9664 3.9242 -2.9263 70.3801 8.6697
MOA Wang et al. (2024b) 27.6122 10.6001 25.4665 5.1210 -2.8837 71.9021 7.8123
MBR Freitag et al. (2023) 27.1221 10.4025 25.3322 4.5642 -2.8912 71.6312 7.9244
PairRank Jiang et al. (2023b) 28.2055 10.8611 25.9361 4.9900 -2.8637 72.0871 7.0420
@@
SpecFuse (w/o Exit) 30.8566 14.0015 28.5648 5.5113 -2.8801 72.8901 4.0983
Table 1: Performance on the English Open-Domain Instruction-Response benchmark, with the best result for each metric highlighted in bold and an $\ast$ indicating the highest result among base LLMs. The upward arrow indicates that a higher value for the metric is better, while the downward arrow indicates that a lower value is better. All ensemble methods in the table integrate all base LLMs, with GF (Qwen2) using the outputs of the other base LLMs as context to generate the fused result through Qwen2-7B. SpecFuse (w/o Exit) refers to not using the Model Exit Mechanism, with all base models continuously participating in the inference process.
Model Bert-S RougeL Avg. Models/Reply
SpecFuse (w/o Exit) 72.8901 28.5648 5.0000
[…]
       w/o Quality Score 72.1122 27.3213 3.2562
       w/o Entropy 72.8105 28.4146 4.1996
Table 2: Ablation study of the Model Exit mechanism. “SpecFuse (w/o Exit)” refers to not utilizing the model exit mechanism. “w/o Quality Score” means that the cumulative quality score is not applied, and models that fail to generate the best segment within the first 10 segments are excluded. “w/o entropy” indicates that entropy is not used to dynamically adjust the softmax temperature coefficient. “Avg. Models/Reply” refers to the average number of models invoked by our framework to respond each question.
Token Genfuse MOA SpecFuse
Num. F.T(s) R.T(s) F.T(s) R.T(s) F.T(s) R.T(s)
100 2.4 5.0 5.1 7.8 0.5 4.2
500 11.5 24.9 25.5 37.7 0.5 19.0
1000 24.3 51.1 51.3 76.9 0.5 37.8
1500 37.2 80.1 80.2 122.9 0.5 58.9
2000 52.9 113.9 114.1 178.0 0.5 80.6
Table 3: Comparison of the time taken by different integration methods for generating varying numbers of tokens. The “F.T” represents First Token Latency, which is the time users wait from inputting data into the system until the system outputs the first token. “R.T” refers to the total time required for the ensemble system to output the complete response to the user.
@@
Gemma-2-9B Gemma et al. (2024) 29.1486 7.6523 18.3456 3.3647 -4.2845$\ast$ 68.7312 9.3586
Qwen2-7B Yang et al. (2024) 29.9296 8.0901 20.0345 3.6181 -4.3271 69.9989 7.1483
Mistral-7B-v0.3 Jiang et al. (2023a) 30.9890$\ast$ 8.6498 20.6568$\ast$ 4.4205 -4.4800 70.1000 7.2251
Glm-4-9B GLM et al. (2024) 30.8761 8.7092$\ast$ 20.4193 4.4674$\ast$ -4.3018 70.2452$\ast$ 5.6893$\ast$
@@
Llama-3-70B AI@Meta (2024) 27.7816 7.0486 20.2227 4.1399 -4.5517 68.5211 8.1475
Qwen2-72B Yang et al. (2024) 31.4356 8.9688 22.4781 4.8838 -4.3368 70.6480 3.6286
Baselines of Ensemble Method
GF (Qwen2) Jiang et al. (2023b) 28.6936 7.8675 18.9339 3.3169 -4.4105 69.8126 9.1109
GF (Mistral) Jiang et al. (2023b) 30.2933 8.1220 20.3324 3.8817 -4.5356 70.0437 7.8613
GF (Glm-4) Jiang et al. (2023b) 30.2643 8.6996 20.5103 4.2722 -4.3316 70.2348 6.1370
MOA Wang et al. (2024b) 30.9631 8.5009 20.5980 4.3586 -4.3100 70.1128 5.9293
MBR Freitag et al. (2023) 30.9335 8.7132 20.6322 4.3149 -4.3060 70.2266 5.3287
@@
Open-Domain Instruction-Response Tasks. We evaluate the performance of our ensemble framework in responding to user queries on both English and Chinese benchmarks and compare it with single LLMs and other ensemble methods. The test results on the English benchmark are shown in Table 1. The experimental results demonstrate that by integrating base LLMs, SpecFuse surpasses all base LLMs and previous ensemble methods across all metrics, with an average improvement of more than $3$ points in the Rouge (n) scores, while also achieving the highest overall ranking in the GPT4-Rank metric, which the responses of all models in the table using GPT-4333https://openai.com/index/gpt-4/. Compared to large models with over $70$B parameters, our method is competitive across most metrics. This suggests that in open-domain scenarios with uneven instruction difficulty, SpecFuse provides more stable output by integrating the advantage of multiple base LLMs, achieving response quality comparable to larger models while significantly lowering deployment costs and complexity. Furthermore, as shown in Table 4, the results on the Chinese benchmark are similar to those on the English benchmark, with SpecFuse outperforming previous ensemble methods and base LLMs across all metrics, demonstrating that its effectiveness is not constrained by language and highlighting its strong generalization ability.
@@
Commonsense, Reasoning, and Instruction-Following Tasks. To investigate the performance of SpecFuse on commonsense, reasoning, and instruction-following tasks, we select three base LLMs with different task specializations and conduct experiments on four benchmarks. As shown in Table 5, SpecFuse (All) outperforms previous ensemble methods across all four benchmarks and is not constrained by the task format. Additionally, by integrating different combinations of base LLMs with SpecFuse, we observe the SpecFuse delivers the most significant overall improvement, with SpecFuse (Qwen2+GLM4) achieving gains of $+2.5$ on MMLU and $+2.4$ on ARC-C by leveraging the respective strengths of Qwen2-7B and GLM-4-9B.
3.3 Effect of Model Exit Mechanism
To analyze the effectiveness of the Model Exit Mechanism, we conduct experiments on the Instruction-Response English benchmark test set using five base LLMs. The experimental results are presented in Table 2. It is evident that SpecFuse with the Model Exit Mechanism reduces the average number of models invoked per response by 50% compared to SpecFuse (w/o Exit), significantly lowering computational resource consumption. Please refer to Appendix C.1 for more experiments and analysis on the effect of the Model Exit Mechanism.
In addition, we designed an ablation study of the Model Exit Mechanism to investigate the roles of Quality Score and the use of Entropy to dynamically adjust the temperature coefficient. The experimental results are shown in Table 4. It is evident that both components have a significant impact on the Model Exit Mechanism, and their combination yields better performance. A more detailed ablation study is provided in Appendix C.2.
3.4 Analysis of Time Consumption
To compare the first-token latency and overall inference time of our framework with previous work, we conduct comparative experiments of different ensemble methods on the Open-Domain IR English benchmark with an integration of 4 base LLMs. Detailed experimental settings can be found in Appendix C.3. The experimental results are shown in Table 3. It can be seen that the First Token Latency of the Genfuse and MOA methods increases as the token generation length grows. When the token generation length reaches $2000$, the MOA method requires the user to wait $114$ seconds to see the output. In contrast, our SpecFuse method outputs in real-time with each generated segment, requiring only $0.5$s for the user to start seeing the output, and this time does not increase with the overall generation length. Additionally, in terms of total inference time, our approach also outperforms the first two complete response fusion methods.
@@ 3.4 Analysis of Time Consumption
Figure 3: The variation in SpecFuse’s RougeL score as the number of base LLMs increases. +Qwen2-7B indicates adding a Qwen2-7B model to the ensemble.
Figure 4: The variation trends of SpecFuse’s BertScore and RougeL score as the maximum generation length of each candidate segment changes.
3.5 Analysis of the Number of Base LLMs
We test the variation in SpecFuse’s performance as the number of base LLMs increases on the test set of the Open-Domain Instruction-Response English benchmark. As shown in Figure 3, the performance of SpecFuse consistently improves as the number of integrated base LLMs increases. When stronger models are introduced, the improvements are substantial, while adding weaker models results in moderate enhancements. This reveals the effect of adding higher-performing models to the system, highlighting the advantage of our framework in seamlessly integrating new LLMs without the need for any training or adjustments.
3.6 Analysis of the Maximum Length of Candidate Segments
To explore the impact of different maximum generation lengths of candidate segments on the performance of the SpecFuse framework, we conducted tests on the English open-domain IR development set. As shown in Figure 4, both the BertScore and RougeL score initially rise with increasing maximum length, reaching their highest point at a length of $10$, after which they begin to decline. The reason for this phenomenon is that when the candidate segment length is too short, it contains insufficient information, which affects the judgment of the verification component and makes it difficult for the models to effectively inspire one another. On the other hand, if the length is too long, the reduced frequency of cross-model interaction leads to less effective knowledge fusion, ultimately diminishing the quality of the final output.
3.7 Case Study
Table 6 presents a case from the SpecFuse workflow. Please refer to Appendix D for details.
@@ 4 Related Work
Model ensemble Sagi and Rokach (2018); Yang et al. (2023) is a widely used technique that enhances forecasting accuracy and robustness by combining predictions from multiple models. With the widespread application of large language models Jiang et al. (2023a), model ensemble has become a hot topic Yu et al. (2023) in LLM research.
Due to limitations such as differences in vocabulary across models, earlier large model ensemble methods Freitag et al. (2023) involved developing ranking models (Salazar et al., 2020; Liu and Liu, 2021) to compare complete candidate results generated by multiple LLMs, selecting the top-ranked candidate as the ensemble output. Subsequently, Lu et al. (2023) first tests base models on a dataset, maps each question to its best-performing model, and trains a routing model that scores the given query and directs it to the highest-scoring model, inferring from that model alone. However, these methods Wang et al. (2024a); Shnitzer et al. (2023) are coarse-grained ensembles that output a complete response generated by a single model, without integrating the strengths of each model, which can limit the potential for producing better outputs.
To address the above issues, Jiang et al. (2023b) uses a trained fusion model that takes the complete responses outputted by the base models as context to generate a fused output. Similarly, Lv et al. (2024) propose the unified ensemble method that first ranks responses, selects the top few candidates as context, and then generates a new response. Subsequently, Huang et al. (2024) first utilize a specific dataset to adapt the relative space and then aggregate the responses. To avoid using additional data for training or adapting models, Wang et al. (2024b) propose the MOA method, which applies multi-layer fusion and selects an Aggregator from the base models to produce the final result.
However, different Aggregators excel at different problems, and using a single Aggregator to output the final result can lead to performance instability in open-domain questions. In addition, these methods integrate complete responses, neglecting the potential for LLMs to collaboratively improve output quality and causing long delays before the fused results begin to be output. In contrast, our SpecFuse method iteratively produces the next segment through collaboration among LLMs during inference, effectively addressing the above issues.